OpenStudy (anonymous):

The areas of two circles are 64 π and 100 π. Find the ratio of their radii and the ratio of their circumference. (Show the solution. Thanks)

OpenStudy (anonymous):

 AREA OF CIRCLE

OpenStudy (campbell_st):

divide both sides by pi 64:100 take the square root of both sides 8:10 so the ratio of the radii is 4:5

OpenStudy (anonymous):

64pi = pi*r^2 64 = r^2 r = 8 in second case R = 10 now 2*pi*r/2*pi*R = 8/10 :D

OpenStudy (saifoo.khan):

Use; Area = pi*r^2 Insert the values of areas.. Then you will get the respective values of "r". Then use then to find ratios. @ellabang